# Enabling Fast 2-bit LLM on GPUs: Memory Alignment, Sparse Outlier, and Asynchronous Dequantization

### 28 Nov 2023 v1

@@ Abstract.
Large language models (LLMs) have demonstrated impressive abilities in various domains while the inference cost is expensive. Many previous studies exploit quantization methods to reduce LLM inference cost by reducing storage and accelerating computation. The state-of-the-art methods use 2-bit quantization for mainstream LLMs (e.g. Llama2-7b, etc.). However, challenges still exist in reducing LLM inference cost with 2-bit quantization: (1) Nonnegligible accuracy loss for 2-bit quantization. Weights are quantized by groups, while the ranges of weights are large in some groups, resulting in large quantization errors and nonnegligible accuracy loss (e.g. ¿3% for Llama2-7b with 2-bit quantization in GPTQ and Greenbit). (2) Limited accuracy improvement by adding 4-bit weights. Increasing 10% extra average bit more 4-bit weights only leads to ¡0.5% accuracy improvement on a quantized Llama2-7b model. (3) Time-consuming dequantization operations on GPUs. Mainstream methods require a dequantization operation to perform computation on the quantized weights, and the 2-order dequantization operation is applied because scales of groups are also quantized. These dequantization operations lead to ¿50% execution time, hindering the potential of reducing LLM inference cost.
@@
To tackle these challenges and enable fast and low-cost LLM inference on GPUs, we propose the following techniques in this paper. (1) Range-aware quantization with memory alignment. We point out that the range of weights by groups varies. Thus, we only quantize a small fraction of groups with the larger range using 4-bit with memory alignment consideration on GPUs. (2) Accuracy-aware sparse outlier. We point out that the distribution of the sparse outliers with larger weights is different in 2-bit and 4-bit groups, and only a small fraction of outliers require 16-bit quantization. Such design leads to ¿0.5% accuracy improvement with ¡3% average increased bit for Llama2-7b. (3) Asynchronous dequantization. We point out that calculating the scales of each group is independent of the loading weights of each group. Thus, we design the asynchronous dequantization on GPUs, leading to up to 3.92 speedup. We conduct extensive experiments on different model families and model sizes. We achieve 2.85-bit for each weight considering all scales/zeros for different models. The end-to-end speedup for Llama2-7b is 1.74 over the original model, and we reduce both runtime cost and hardware cost by up to 2.70 and 2.81 with less GPU requirements.
1. Introduction
@@
Large language models (LLMs) demonstrate remarkable capabilities in various domains, excelling in tasks like natural language understanding and generation (Du et al., 2022; Ni et al., 2023). However, their computational expense during inference is noteworthy. For instance, OpenAI reveals that the GPT-4 Turbo incurs a cost of millions of dollars each day (Patel and Ahmad, 2023). This substantial cost underscores the financial considerations associated with deploying and utilizing such advanced language models (Peng et al., 2023). Many previous studies exploit quantization methods to reduce LLM inference cost by reducing storage and accelerating computation (Frantar et al., 2022; Lin et al., 2023; Dettmers et al., 2023; Kim et al., 2023; Xiao et al., 2023). The state-of-the-art methods have realized 2-bit quantization for mainstream LLMs, exemplified by the Llama-2 family (Touvron et al., 2023). These methods quantize weights by group to improve accuracy and the scales are also quantized (2-order) to further reduce average bit (Guo et al., 2023).
@@
However, there are still existing challenges of reducing LLM inference cost with the 2-bit quantization method. (1) Nonnegligible accuracy loss for 2-bit quantization. Both the post-training quantization (PTQ, e.g., GPTQ (Frantar et al., 2022)) and the quantization-aware training (QAT, e.g., LLM-QAT (Liu et al., 2023) and Greenbit (Guo et al., 2023)) approaches incur the 3.2% to 5.6% accuracy loss for Llama2-7b with 2-bit quantization (Frantar et al., 2022; Guo et al., 2023). (2) Limited accuracy improvement by adding 4-bit weights. LLMs quantized with 4-bit exhibit an impressive ability to retain accuracy, showcasing less than 1% accuracy loss. Consequently, it is a common practice to improve accuracy by adding 4-bit quantization. However, Figure 2 left middle illustrates that increasing 10% extra average bit with more 4-bit weights only leads to ¡0.5% accuracy improvement on a quantized Llama2-7b model. (3) Time-consuming dequantization operations on GPUs. Mainstream methods require a dequantization operation to perform computation on 2-bit weights. Furthermore, a 2-order dequantization operation is applied because the scales of groups are also quantized. Figure 2 left bottom shows that, the dequantization operation takes over 50% of the total execution time. These challenges become the crucial factors impeding fast and low-cost LLM inference.
@@
Accuracy-aware sparse outlier. We point out that only a small fraction of outliers exist in weights quantized using 2-bit. We quantize these sparse outliers with ¡3% increased average weight bit and improve the accuracy by ¿0.5%.
@@
We conduct extensive experiments on different model families (e.g., Llama2 (Touvron et al., 2023) and ChatGLM3 (Zeng et al., 2022)) and model sizes (e.g., Llama2-7b to Llama2-70b). We achieve 2.85-bit for each weight considering all scales/zeros for different models. The end-to-end speedup for Llama2-7b is 1.74, and we reduce both runtime cost and hardware cost by up to 2.70 and 2.81 with less GPU requirements.
2. Preliminaries and Backgrounds
2.1. Uniform Quantization
A group of weights with range of are quantized to range . Scale () and Zero-point () map the floating point weights to bits integers. Scale is the half-precision step size of the quantizer and Zero-point is an bits integer (Nagel et al., 2021). The quantization and dequantization operations are:
For 2-bit quantization, weights with input dimensions are quantized by group, while the scales of groups within output dimensions are also quantized to further reduce the average bit. Here, the average bit, denoted as with 1-order and 2-order quantization is as follows:
| (5) | 1-order | ||||
| (6) | 2-order |
where represents the size of group and represents the bit width with 2-order.
2.2. Mixed-precision Quantization
In LLMs, certain weights exhibit disproportionately magnitudes compared to the majority. As shown in Figure 3, by applying normal uniform quantization, the quantization range is expanded by these outliers and thus, the majority of weights are quantized with few steps, which leads to large quantization error. Mixed-precision quantization methods like SpQR and SqueezeLLM filter and retain outliers that negatively impact quantization error and store outliers with 16-bit sparse matrix (Dettmers et al., 2023; Kim et al., 2023). These methods quantize the weights matrix into a dense matrix and a 16-bit sparse matrix. The Compressed Sparse Row (CSR) is used to represent sparse matrices, optimizing memory usage. In CSR, three arrays encapsulate the matrix data. The values array (values) efficiently arranges non-zero elements in row-major order, while the column indices (col_ind) array meticulously records the corresponding column index for each non-zero element. The row pointer (row_ptr) array serves as a guide, storing the starting index for each row in the values array. Outlier retention of weights is common for low-bit quantization in LLMs. Each outlier is stored using one 32-bit combination: a 16-bit weight value and a corresponding 16-bit column index. Additionally, for every row, a 32-bit number is allocated to store the total count of outliers up to that row. This results in an overall 32-bit storage overheads for each weight outlier.
3. Range-aware Quantization with Memory Alignment
@@
Challenge. The quantization method is widely used to reduce LLM inference cost (Frantar et al., 2022; Lin et al., 2023; Dettmers et al., 2023) by reducing storage and accelerating computation. Quantization uses discrete values to represent continuous weights, leading to quantization errors. We depict the distribution of weights in typical LLMs and the comparison of using 2-bit and 4-bit quantization in Figure 5(a). We find that the 2-bit approach leads to larger quantization errors and further results in larger accuracy loss. Thus, the key challenge is that even the state-of-the-art 2-bit methods still suffer from ¿3% accuracy loss for Llama2-7b (Frantar et al., 2022; Guo et al., 2023).
@@
Motivations and Insights. Some previous methods like SpQR (Dettmers et al., 2023), SparseGPT (Frantar and Alistarh, 2023) and AWQ (Lin et al., 2023) have analyzed that the quantization errors are large in specific groups, so we further depict the range of weights per group, as shown in Figure 5(b). The range of weights exhibits different distributions for different groups, and only a small fraction of groups show a large range while the others show a small range. Applying 4-bit quantization to these groups with a large range can reduce the quantization error. Thus, our key insight is, only a small fraction of weights requires 4-bit quantization and average bit can be still reduced by applying 2-bit quantization on the other groups.
Approach: Range-aware Quantization. In Figure 4, we depict the quantization flow of weights. Notably, weights are quantized by group to reduce accuracy loss while the groups with a large range still require larger bit-width to improve accuracy. Therefore, we first analyze the range distribution by group. The range of groups in the same input channels is merged to streamline the complexity of hardware computation. Concretely, weights are divided by the Hessian inverse matrix to get the calibrated weights, leveraging the Hessian inverse matrix’s ability to discern weight importance. Subsequently, the calibrated weights are squared to magnify the range variance. The amplitude of each input channel is computed as follows:
@@
Thus, the range of weights by groups is equivalent to the amplitude variance in the grouped input channels. Then, we analyze the variance and obtain the information that the input channels with large variance require 4-bit quantization and others require 2-bit quantization. According to the analysis results, we apply the QAT finetuning approach to reduce the quantization error. In the forward process of QAT, the weights with a large range are quantized and then dequantized with 4-bit by group while others are with 2-bit. After finetuning, we use the same analysis results to quantize different grouped input channels with different bit-width. For small range input channels, we first do 2-bit 1-order quantization for the weights by group in the direction of input channels to get the quantized weight , 1-order zeros and scales . To further reduce the average bit, the scales are also quantized with 4-bit by group in the direction of output channels (2-order quantization) to get 1-order quantized scales , 2-order zeros and scales . During 2-bit quantization, outliers are also detected and retained to reduce accuracy loss. Then, the weights in large range input channels are quantized with 4-bit.
@@
Approach: Memory Alignment. As illustrated in Figure 4, the quantized weights are split into two quantized weight matrixes. The first matrix stores 3 groups of 2-bit quantized weights (the size of group is 16) and 8 4-bit quantized weights into continuous memory. The other matrix stores 8 remaining 4-bit quantized weights into independent memory. Thus, the memory access for each thread in GPU is memory-aligned. Then, 1-order zeros and scales are packed together and two of three scales are compressed into two 3-bit to fill 16-bit memory without paddings while only the 2-order zeros are stored with paddings. These memory alignment methods preserve accuracy loss and only leads to ¡0.02 average bit overheads.
4. Accuracy-aware Sparse Outlier
Challenge. The mixed-precision quantization method can efficiently reduce quantization errors of channels with large variations. However, there are still some outliers sparsely distributed in both 2-bit and 4-bit channels. Outliers in 2-bit channels lead to significant quantization errors. In order to reduce the quantization error of these outliers, one straightforward approach is to continue adding 4-bit channels with the increased average bit. Nevertheless, the accuracy improvement of adding 4-bit channels is limited because these outliers are sparsely distributed. Figure 2 left middle shows that, increasing ¿10% extra bit only leads to ¡0.5% accuracy improvement when adding more 4-bit channels. Thus, the key challenge lies in reducing the quantization error of these sparse outliers with less increased average bit.
Motivation and Insights. Besides straightforwardly adding 4-bit channels, another way to reduce quantization errors is to use the 16-bit weights for these sparse outliers. Previous designs like SpQR (Dettmers et al., 2023) and SqueezeLLM (Kim et al., 2023) introduce 1% 16-bit weights to represent these outliers with the corresponding sparse format(e.g., CSR). Each sparse outlier requires at least one 16-bit for the weight representation and one 16-bit for the position. Thus, introducing 1% 16-bit sparse outliers leads to extra average bit, which still increases ¿10% extra bit. We depict the sparse outlier distribution in 2-bit and 4-bit channels of Llama2-7b in Figure 6 left. Although only 25% channels are quantized using 4-bit in our practice, 70% outliers are distributed in these 4-bit channels. Previous designs using 4-bit quantization (Lin et al., 2023; Frantar et al., 2022) lead to negligible accuracy loss, proving that 4-bit quantization is a promising way for accurate LLM inference. Thus, our key insight is, only a small fraction (e.g., 30%) of sparse outliers are in the 2-bit channels and require 16-bit quantization to maintain the algorithm accuracy.
Approach: Accuracy-aware Sparse Outlier. Based on the mixed 2-bit/4-bit quantization proposed in Sec. 3, we quantize a small fraction (e.g., 25%) of weights using 4-bit by channels. We profile the accuracy improvement by continuing to add 4-bit channels. Figure 6 right shows that, the accuracy improvement is limited by adding ¿25% 4-bit weights. We only consider the average bit of weights, while the scales and zeros are not considered. The actual average bit is also related to the group size. The smaller group size leads to more scales and zeros for each group, resulting in a larger average bit (Guo et al., 2023). In practice, we set the group size to 16, which is larger than Greenbit (Guo et al., 2023) and contributes to a smaller average bit. Instead of adding 4-bit channels when the average bit reaches 2.5, we represent sparse outliers only in 2-bit channels using 16-bit quantization. In practice, we finetune the model and restrict the ratio of sparse outliers in 2-bit channels to 0.2%. Thus, we increase 0.06 average bit, which is less than 3%. For Llama2-7b, the accuracy is increased by 0.6%.
Approach: Sparse Outlier Format and GPU Kernel. We utilize the CSR format used in previous designs (Dettmers et al., 2023; Kim et al., 2023) for sparse outliers representation, and apply GPU kernels in the NVIDIA cuSPARSE library (cuS, [n. d.]) to perform matrix multiplication on these outliers. In practice, only 0.2% weights are quantized using 16-bit, and the execution time on these sparse outliers is much faster (e.g., ¿10 according to our profiling) than the computation on the dense 2-bit/4-bit weights.
5. Asynchronous Dequantization
@@
Challenge. Because the weights are quantized by 2-bit and 4-bit, it requires the dequantization operation to restore the weights to 16-bit before performing the multiplication between the input and weights. We use 1-order and 2-order quantization to scale the weights and scales of each group. Previous designs (e.g., SPQR(Dettmers et al., 2023), LLM-QAT(Liu et al., 2023)) use the synchronous dataflow (i.e., performing dequantization after loading all weights), resulting in ¿50% overheads of end-to-end execution time. Thus, the key challenge is the time-consuming synchronous dequantization operation becomes the bottleneck in accelerating LLM inference after quantization.
@@
Approach: Asynchronous dequantization. Based on the insights above, we design the asynchronous dequantization dataflow as illustrated in Figure 7 on GPUs. With the help of the shared memory, we can overlap the calculating scales of each group and the loading weights of each group. Further, we use CUDA primitive __shfl_down_sync() to reduce the result inside a warp efficiently. From the perspective of memory access, we apply vector load (e.g. double4) technique to load quantized weights, activation, scales and zeros to minimize the numbers of memory access. Algorithm 1 shows the four main parts of our kernel. The preprocess (line 1 3) loads the zeros () and scales ( and ) for the 2-order dequantization, corresponding to ➀ in Figure 7. Then the 2-order dequantization and loading weight and (line 4) provide the scales, zeros and weights for the 1-order dequantization, corresponding to operation ➁. The third part is 1-order dequantization and loading input vector (line 7 8), corresponding to operation ➂. After that, the multiplication (line 10), operation ➃ in Figure 7 is performed. The final part (line 11) is used to reduce data from all threads.
| Method | PIQA/Hell./Wino./ARC-e | Average | |
| Llama2-7b | 16 | 77.2%/56.1%/68.7%/70.2% | 68.0% |
| GPTQ | 3 | 71.7%/48.2%/61.2%/56.1% | 59.3% |
| Greenbit | 2.91 | 77.2%/53.8%/65.8%/62.5% | 64.8% |
| Ours (group=16) | 2.85 | 75.9%/51.3%/66.4%/66.9% | 65.1% |
| Ours (0.2%) | 2.91 | 76.1%/52.0%/66.8%/67.2% | 65.7% |
| Llama2-13b | 16 | 78.8%/59.7%/69.6%/73.3% | 70.4% |
| GPTQ | 3 | 75.2%/56.1%/64.3%/64.2% | 64.0% |
| Ours (group=16) | 2.85 | 76.9%/55.6%/66.2%/69.2% | 67.0% |
| Ours (0.2%) | 2.91 | 77.3%/55.8%/66.7%/69.5% | 67.6% |
| Llama2-70b | 16 | 81.1%/64.0%/77.0%/77.7% | 75.0% |
| Ours (group=8) | 2.91 | 80.5%/62.1%/74.9%/76.5% | 73.5% |
| ChatGLM3-6b | 16 | 70.8%/49.4%/61.3%/51.1% | 58.1% |
| Ours (group=8) | 2.91 | 66.7%/44.7%/59.4%/48.5% | 54.8% |
@@
| Our | A1001 | 400/400 | 50.5 | 1.31 | 2.62 | 2.42 | |
| PyTorch | 30901 | 350/350 | 32.5 | 1 | 1 | 1 | |
| ChatGLM3-6b | Our | 30901 | 350/350 | 40.0 | 1.23 | 1.23 | 1.23 |
| Our | 20801 | 170/215 | 45.5 | 1.40 | 2.70 | 2.75 |
6. EXPERIMENTAL RESULTS
6.1. Experimental Setup
Benchmarks. We conduct comprehensive experiments on the Llama2-7b model family (Touvron et al., 2023) and the ChatGLM3-6b model (Zeng et al., 2022). The accuracy of the introduced quantized LLMs is assessed across four zero-shot benchmarks, namely PIQA (Bisk et al., 2020), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC-e (Clark et al., 2018). We use 2-order weight-only quantization to evaluate the performance.
Baselines. We compare our design with state-of-the-art quantization designs, including GPTQ (Frantar et al., 2022) and Greenbit (Guo et al., 2023). The original PyTorch implementation on HuggingFace (Hug, [n. d.]) is used as the baseline.
Hardware Platforms. We implement our design and compare other baselines on NVIDIA RTX 2080, NVIDIA RTX 3090, and NVIDIA A100 (80G) GPUs with CUDA version 12.2. The runtime cost is evaluated considering the power consumption using the NVML library (nvm, [n. d.]), and the hardware cost is evaluated considering the price ($12,500 for A100, $1499 for RTX 3090, and $699 for RTX 2080) with the 5-year service life period.
6.2. Accuracy Evaluations
Table 1 shows algorithm accuracy with zero-shot performance on Llama2 family and ChatGLM3-6b. We set the group size to 16 for Llama2-7b and Llama2-13b with 0.2% sparse outliers, and the group size to 8 for Llama2-70b and ChatGLM3-6b.
Compared with the state-of-the-art GPTQ (Frantar et al., 2022) method, our design significantly reduces the accuracy loss from 8.7% to 2.3% for Llama2-7b, and from 6.4% to 2.8% Llama2-13b. Compared with Greenbit (Guo et al., 2023) on Llama2-7b, we achieve 0.9% higher accuracy with the same 2.91 average bit. On Llama2-70b and ChatGLM3-6b, the accuracy loss is also controllable from 1.5% to 3.3%.
6.3. Performance Evaluations
We compare the inference cost (end-to-end speedup, runtime cost, and hardware cost) among our method, PyTorch, and GPTQ in various model families and model sizes in Table 2. All results are normalized to PyTorch. In the Llama2 model family, our design achieves end-to-end speedup and far outperforms other previous works (e.g., GPTQ(Frantar et al., 2022)) in terms of running cost reduction, reaching over the original model. Moreover, the hardware requirement is dramatically lowered in our design for all models in Table 2, bringing up to runtime cost reduction and hardware cost reduction. The performance underscores the effectiveness of our design’s optimizations and the practical utility of our approach tailored to optimize the inference and deployment of LLMs.
6.4. Detailed Discussions
6.4.1. Kernel Speedup
6.4.2. Integrate with FlashAttention
FlashAttention (Dao et al., 2022) is widely used to accelerate LLM inference. We show that our method is also compatible with FlashAttention. Figure 9 shows that our method achieves higher speedup compared with adopting FlashAttention. Integrating our method with FlashAttention further accelerates LLM inference by up to , and solves the limited memory problem on GPUs like NVIDIA RTX 2080.
7. Conclusions
We enable fast and low-cost LLM inference on GPUs in this paper with three novel techniques. We quantize 25% weights 4-bit quantization and 0.2% sparse outliers with 16-bit based on insights into weight distribution in LLMs. We further design the asynchronous GPU kernel to accelerate LLM dequantization. For Llama2-7b, our method achieves 2.85 average bits, and the end-to-end speedup is 1.74. We reduce both runtime cost and hardware cost by up to 2.70 and 2.81 with less GPU requirements.

### 13 Dec 2023 v2

# Enabling Fast 2-bit LLM on GPUs: Memory Alignment and Asynchronous Dequantization

@@ Abstract
Large language models (LLMs) have demonstrated impressive abilities in various domains while the inference cost is expensive. Many previous methods use 2-bit quantization to reduce LLM inference cost by reducing storage and accelerating computation. However, challenges still exist in reducing LLM inference cost: (1) Large accuracy loss for 2-bit quantization. Weights are quantized by groups, while the gradient of weight are large in some groups, resulting in large quantization errors and accuracy loss. (2) Time-consuming dequantization operations on GPUs. Mainstream methods require a dequantization operation to perform computation on the quantized weights, and the 2-order dequantization operation is applied because scales of groups are also quantized. These dequantization operations lead to >50% execution time, hindering the potential of reducing LLM inference cost.
@@
To tackle these challenges, we propose the following techniques in this paper. (1) Range-aware quantization with memory alignment. We only quantize a small fraction of groups with the larger range using 4-bit with memory alignment consideration on GPUs. (2) Asynchronous dequantization. We point out that calculating the scales of each group is independent of the loading weights of each group. Thus, we design the asynchronous dequantization on GPUs, leading to up to 3.92 speedup. Our method achieves 2.85-bit for each weight considering all scales/zeros for different models. The end-to-end speedup for Llama2-7b is 1.74 over the original model.
1 Introduction
@@
Large language models (LLMs) demonstrate remarkable capabilities in various domains like natural language understanding and generation [1, 2]. However, their computational expense during inference is noteworthy. For example, OpenAI reveals that the GPT-4 Turbo incurs a cost of millions of dollars each day [3][4]. Many previous studies exploit quantization methods to reduce LLM inference cost by reducing storage and accelerating computation [5, 6, 7, 8, 9]. The state-of-the-art methods have realized 2-bit quantization for mainstream LLMs, exemplified by the Llama-2 family [10]. These methods quantize weights by group to improve accuracy and the scales are also quantized (2-order) to further reduce average bit [11].
@@
However, there are still existing challenges of 2-bit quantization. (1) Large accuracy loss for 2-bit quantization. Both the post-training quantization (PTQ, e.g., GPTQ [5]) and the quantization-aware training (QAT, e.g., LLM-QAT [12] and Greenbit [11]) approaches incur the 3.2% to 5.6% accuracy loss for Llama2-7b with 2-bit quantization [5, 11]. (2) Time-consuming dequantization operations on GPUs. Mainstream methods require a dequantization operation to perform computation on 2-bit weights. Furthermore, a 2-order dequantization operation is applied because the scales of groups are also quantized. The dequantization operation takes over 50% of the total execution time. These challenges become the crucial factors impeding fast and low-cost LLM inference.
@@
We conduct extensive experiments on Llama-2 model families. We achieve 2.85-bit for each weight considering all scales/zeros for different models. The end-to-end speedup for Llama2-7b is 1.74, and we reduce both runtime cost and hardware cost by up to 2.70 and 2.81 with less GPU requirements.
2 Method
2.1 Range-aware Quantization
@@
Challenge. Quantization uses discrete values to represent continuous weights, leading to quantization errors. Applying 2-bit quantization leads to larger quantization errors and further results in larger accuracy loss for Llama2-7b. Even the state-of-the-art 2-bit methods still suffer from >3% accuracy loss for Llama2-7b [5, 11].
@@
Motivations and Insights. Some previous methods like SpQR [7] and AWQ [6] have analyzed that the quantization errors are large in specific groups. Applying 4-bit quantization to these groups with a large range can reduce the quantization error. Thus, our key insight is, only a small fraction of weights requires 4-bit quantization and average bit can be still reduced by applying 2-bit quantization on the other groups.
@@
Approach: 2/4-bit Quantization. In Figure 1, we depict the quantization flow of weights. First, we analyze the range distribution by group. The range of groups in the same input channels is merged to streamline the complexity of hardware computation. Thus, the range of weights by groups is equivalent to the amplitude variance in the grouped input channels. Then, we analyze the variance and obtain the mixed-precision layout that the input channels with large variance require 4-bit quantization and others require 2-bit quantization. According to the mixed-precision layout, we apply the finetuning approach and then quantize different grouped input channels with different bit-width. For small range input channels, we quantize the weights with 2-bit and also quantize the scales with 4-bit to reduce average bit. For large range input channels, we quantize the weights with 4-bit. Equation 1 is the average bit of 2/4-bit quantization.
where is proportion of weights quantized with 2-bit, and are the group sizes of 1-order and 2-order quantization separately, is the bit precision of 2-order quantization.
@@
Approach: Memory Alignment. As illustrated in Figure 1, the quantized weights are split into two quantized weight matrixes. The first matrix stores 3 groups of 2-bit quantized weights (the size of group is 16) and 8 4-bit quantized weights into continuous memory. The other matrix stores 8 remaining 4-bit quantized weights into independent memory. Thus, the memory access for each thread in GPU is memory-aligned. Then, the zeros and scales are packed together and two of three scales are compressed to fill 16-bit memory. These memory alignment approaches preserve accuracy loss and only lead to <0.02 average bit overheads.
2.2 Asynchronous Dequantization
@@
Challenge. Before the multiplication operations, the quantized weights are required to be dequantized to 16-bit. Previous designs (e.g., SPQR[7]) use the synchronous dataflow (i.e., performing dequantization after loading all quantized weights, scales and zeros), resulting in >50% overheads of end-to-end execution time. Thus, the key challenge is the time-consuming synchronous dequantization operation becomes the bottleneck in accelerating LLM inference after quantization.
@@
Approach: Asynchronous dequantization. Based on the insights above, we design the asynchronous dequantization dataflow as illustrated in Figure 2 on GPUs. With the help of the shared memory, we can overlap the calculating scales of each group and the loading weights of each group. Further, we use CUDA primitive __shfl_down_sync() to reduce the result inside a warp efficiently. From the perspective of memory access, we apply vector load (e.g. double4) technique to load quantized weights, activations, scales and zeros to minimize the numbers of memory access.
3 Experiments
We conduct comprehensive experiments on the Llama2-7b model family [10]. The accuracy of the introduced quantized LLMs is assessed across four zero-shot benchmarks, namely PIQA [13], HellaSwag [14], WinoGrande [15], ARC-Easy [16]. We use 2-order weight-only quantization to evaluate the performance.
Baselines. We compare our design with state-of-the-art quantization designs, including GPTQ [5] and Greenbit [11]. The original PyTorch implementation on HuggingFace [17] is used as the baseline.
Hardware Platforms. We implement our design and compare other baselines on NVIDIA RTX 2080, NVIDIA RTX 3090, and NVIDIA A100 (80G) GPUs with CUDA version 12.2. The runtime cost is evaluated considering the power consumption using the NVML library [18], and the hardware cost is evaluated considering the price ($12,500 for A100, $1499 for RTX 3090, and $699 for RTX 2080) with the 5-year service life period.
| Method | PIQA | HellaSwag | WinoGrande | ARC-Easy | Average | |
| Llama2-7b | 16 | 77.2% | 56.1% | 68.7% | 70.2% | 68.0% |
| GPTQ | 3 | 71.7% | 48.2% | 61.2% | 56.1% | 59.3% |
| Greenbit | 2.91 | 77.2% | 53.8% | 65.8% | 62.5% | 64.8% |
| Ours | 2.85 | 75.9% | 51.3% | 66.4% | 66.9% | 65.1% |
| Llama2-13b | 16 | 78.8% | 59.7% | 69.6% | 73.3% | 70.4% |
| GPTQ | 3 | 75.2% | 56.1% | 64.3% | 64.2% | 64.0% |
| Ours | 2.85 | 76.9% | 55.6% | 66.2% | 69.2% | 67.0% |
@@ 3 Experiments
| Our | A1001 | 400/400 | 50.5 | 1.31 | 2.62 | 2.62 |
Accuracy Evaluations. Table 1 shows algorithm accuracy with zero-shot performance on Llama2 family. We set the group size to 16 for Llama2-7b and Llama2-13b. Compared with the state-of-the-art GPTQ [5] method, our design significantly reduces the accuracy loss from 8.7% to 2.9% for Llama2-7b, and from 6.4% to 3.4% Llama2-13b. Compared with Greenbit [11] on Llama2-7b, we achieve 0.3% higher accuracy with the lower average bit.
Kernel Speedup. We compare our kernel performance with some mainstream quantized kernels(e.g., GPTQ [5]) and the original FP16 models in common LLM GEMV cases. Our kernel achieves over FP16 on various NVIDIA GPUs.
Performance Evaluations. We compare the inference cost (end-to-end speedup, runtime cost, and hardware cost) among our method, PyTorch, and GPTQ in various model families and model sizes in Table 2. All results are normalized to PyTorch. In the Llama2 model family, our design achieves end-to-end speedup and far outperforms other previous works (e.g., GPTQ[5]) in terms of running cost reduction, reaching over the original model. Moreover, the hardware requirement is dramatically lowered in our design for all models in Table 2, bringing up to runtime cost reduction and hardware cost reduction. The performance underscores the effectiveness of our design’s optimizations and the practical utility of our approach tailored to optimize the inference and deployment of LLMs.